The rms velocity of molecule in a sample of helium gas is 5/7th of the molecules in the sample of hydrogen If the temperature of hydrogen samples is 0 degree celcius then the temperature of helium sample about

Grade:12

1 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallvi

 we know

Vrms = √3RT/M

Vrms(H)  =√3R273/2

Vrms(He) = √3RT/4

 

Vrms(He) =5/7 Vrms(H)

√3RT/4  =5/7 { √3R273/2 }

T/4  = 25/49 {273/2}

T=278.57 K

  = 5.57 C

Please feel free to post as many doubts on our discussion forum as you can.
If you find any question Difficult to understand - post it here and we will get you the answer and detailed
solution very  quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin